### v1

@@ Abstract
This paper develops a theory of bargaining in firm-to-firm trade with two-sided market power. The framework accommodates flexible market structures, yielding analytical expressions for pair-specific markups and pass-through elasticities. In U.S. import data, we estimate strong importer bargaining power and an upward-sloping export supply curve, consistent with oligopsony power. Pass-through of the 2018 tariffs in firm-to-firm relationships is incomplete, in contrast to product-level studies, primarily due to exporter cost reductions driven by falling demand from dominant buyers. Our study highlights how bargaining and network rigidities shape price outcomes, with implications for markup dispersion and shock propagation in global value chains.
@@ 1 Introduction
The recent resurgence of protectionist trade policies has reignited debate over how tariffs affect international prices. This question is particularly salient in today’s global economy, where production is fragmented across global value chains and most trade occurs within ongoing relationships between importers and exporters. These firm-to-firm linkages shape the cross-border transmission of shocks (Boehm et al., 2019); however, their implications for price setting and tariff pass-through remain less well understood.
A growing body of literature highlights the limitations of standard trade models in this context. Rather than engaging in anonymous spot-market transactions, firms often trade within long-term relationships shaped by contract incompleteness, relationship-specific investments, and lock-in frictions (Antràs, 2015). In this environment, prices arise from bilateral negotiations between exporters and importers, both of whom may hold significant market power (Antràs and Staiger, 2012; Morlacco, 2019; Dhyne et al., 2022).
This paper develops a theory of bargaining in firm-to-firm trade that incorporates two-sided market power and network rigidities. The framework accommodates flexible upstream and downstream market structures and delivers closed-form expressions for markups and pass-through elasticities as functions of market shares and a small set of structural parameters. We estimate these parameters using U.S. import data, validate the model’s cross-sectional predictions against observed pricing patterns, and use the estimated framework to quantify aggregate pass-through in firm-to-firm trade, highlighting how bargaining and network structure shape both the level and sources of pass-through in response to trade shocks.
Section 2 introduces the theory. We consider a fixed network of importer-exporter (or buyer-supplier) relationships, where each pair negotiates bilaterally over the price of an intermediate input. Given the price, the importer unilaterally chooses the traded quantity to minimize costs, making prices allocative. In the event of a disagreement, each party falls back on trade with its other existing partners. As a result, the network structure influences the value of each party’s outside option, and in turn, their effective bargaining power.
Our framework offers a unified theory of exporter and importer market power in firm-to-firm trade. Exporter rents arise from three sources: monopoly power through product differentiation, oligopoly power due to supplier concentration, and quasi-rents from an upward-sloping residual supply curve. When exporters hold full bargaining power, the model nests standard trade settings in which prices reflect an oligopoly markup over marginal cost (Atkeson and Burstein, 2008; Dhyne et al., 2022). This markup increases with the exporter’s share of the buyer’s total input purchases—a measure we refer to as the exporter’s supplier share.
Our first contribution is to characterize the role of importer market power in price-setting. This power arises when suppliers earn economic rents, creating scope for buyers to extract surplus by negotiating lower prices. It also reflects oligopsony power, as concentrated buyers can negotiate lower prices. When importers hold full bargaining power, prices can fall below marginal cost, generating a negative price–cost margin akin to an oligopsony markdown. This markdown declines with the importer’s share of the supplier’s total output–a measure we refer to as the importer’s buyer share. [^1] Our definition of the importer’s markdown differs from the standard price–marginal revenue wedge. See Section 2 for details. In general, the negotiated price reflects a convex combination of oligopoly and oligopsony forces, with weights determined by the firms’ outside options and other exogenous determinants of bargaining power.
Our second contribution is to characterize how importer market power shapes the pass-through of tariff shocks to bilateral prices. We focus on short-run pass-through elasticities, derived as local perturbations around the static price equilibrium, while holding the network and general equilibrium conditions fixed. This isolates the intensive margin of price adjustment from longer-run responses driven by entry, reallocation, or macroeconomic feedback.
Importer market power affects pass-through through two distinct mechanisms: a markup channel and a cost channel. The markup channel captures how bilateral markups respond to price changes. When exporters hold full bargaining power, markups exhibit strategic complementarities, a well-known source of incomplete pass-through (Amiti et al., 2014, 2019a). With two-sided market power, the markup elasticity becomes a convex combination of strategic complementarities and strategic substitutabilities, as importers reduce markdowns in response to higher prices, amplifying pass-through.
The cost channel captures how tariffs affect the exporter’s marginal cost through scale effects. While prior work links pass-through to cost adjustment under decreasing returns to scale (e.g., Burstein and Gopinath 2015), our framework introduces a new source of heterogeneity: the elasticity of residual supply increases with an importer’s buyer share. As a result, larger buyers face more incomplete pass-through. Although both channels interact to shape price responses, the cost channel dominates when bargaining power is two-sided and returns to scale are decreasing. In this setting, markup adjustments are modest and offset by opposing strategic forces, while cost-side responses amplify the heterogeneity of pass-through.
Section 3 describes the data and presents preliminary empirical evidence. Our main data source is the U.S. Census Longitudinal Firm Trade Transaction Database (LFTTD), which reports prices and quantities for each import transaction involving a U.S. importer, along with identifiers for the foreign exporter and the 10-digit product code. These features allow us to construct bilateral market shares on both sides of the transaction, which are the key sufficient statistics in the model governing the distribution of markups and pass-through elasticities. Our main sample focuses on repeated, arm’s-length relationships involving the exchange of intermediate inputs, excluding related-party links as flagged in the LFTTD.
Pricing and pass-through patterns in the data closely align with the model’s predictions under two-sided market power. Controlling for supplier–product–time fixed effects to proxy for marginal costs, we find that prices rise with the exporter’s supplier share and fall with the importer’s buyer share, consistent with the presence of both oligopoly and oligopsony forces. In contrast, tariff pass-through declines with buyer share and is non-decreasing in supplier share, suggesting that price adjustments are primarily driven by the cost channel, with a more limited role for the markup channel, as expected under two-sided market power.
These empirical comovements reinforce the model’s central mechanisms and are difficult to reconcile with conventional pricing frameworks that assume one-sided or purely competitive behavior. Nonetheless, reduced-form estimates may conflate the effects of competitor behavior, input price variation, or broader macroeconomic shocks, limiting their interpretability. A structural approach is therefore needed to isolate the underlying mechanisms. [^2] For a related discussion in a different context, see Berger et al. (2022).
To do so, Section 4 develops a structural estimation strategy around two key parameters: the importer’s bargaining power and the exporter’s returns to scale, which determines the slope of the input supply curve. We estimate that U.S. importers hold bargaining power of approximately 0.8, roughly four times greater than that of their foreign suppliers, and that the returns to scale parameter is about 0.45, consistent with upward-sloping supply. Identification exploits cross-sectional price variation across U.S. importers within a supplier–product–year. As the estimating equations hold independently of the general equilibrium environment, the approach remains valid without explicitly modeling broader equilibrium feedback.
We assess the model’s validity by evaluating its ability to replicate both the average and heterogeneous effects of the 2018 tariffs on prices and quantities. First, the model closely matches the average price response as well as its variation with buyer and supplier shares, consistent with the role of oligopoly, oligopsony, and cost-side adjustments. Second, applying the IV-based test proposed by Adão et al. (2023), we find that our model outperforms alternatives that either impose constant returns to scale or exclude bargaining. Third, while the model underestimates the magnitude of quantity changes, it successfully captures their direction, lending support to its core allocative mechanism. Together, these results validate the model’s key assumptions and suggest that general equilibrium feedback or omitted strategic interactions are unlikely to materially bias its pass-through predictions.
In Section 5, we use the estimated model to assess how the 2018 Trump tariffs affected aggregate U.S. import prices through firm-to-firm relationships. Rather than providing a full accounting of tariff incidence, which lies beyond the scope of our framework, we focus on one key margin: price adjustment within ongoing matches, and its decomposition into changes in markups and marginal costs.
We find that aggregate pass-through in firm-to-firm trade is substantially incomplete, with elasticities ranging from 65 to 71%. The model sheds light on the mechanisms behind this muted response. The key insight is that most of the adjustment occurs on the cost side: exporters facing weaker demand from powerful buyers move down their marginal cost curves, absorbing a sizable share of the tariff burden. Markup changes occur but contribute little to the aggregate response due to offsetting effects of strategic complementarities and substitutabilities. This distinction is important. While prior work often attributes incomplete pass-through to variable markups and strategic pricing, our results highlight the dominant role of cost adjustment and supply-side forces.
Our results stand in contrast to studies documenting near-complete pass-through of the 2018 tariffs (e.g., Amiti et al., 2019b, 2020; Cavallo et al., 2020; Fajgelbaum et al., 2020; Flaaen et al., 2020). We reconcile this divergence on two grounds. First, our analysis focuses on intermediate inputs and arm’s-length transactions, where pass-through may be lower due to input specificity, contractual frictions, or capacity constraints. Second, we isolate pricing within repeated matches between importers and exporters. By contrast, product-level unit values used in aggregate studies conflate within-match price changes with compositional shifts in trading partners, notably including one-off exchanges. Such one-off transactions are more likely to reflect marginal cost pricing and exhibit full pass-through. [^3] See Cajal-Grossi et al. (2023) for evidence that markups are lower and prices more competitive in spot transactions than in relational matches.
@@
Our paper contributes to the literature on pricing and shock transmission in firm-to-firm trade, with a particular focus on the role of importer market power. A growing body of empirical work highlights the influence of dominant buyers in shaping supplier outcomes. Bernard et al. (2019) and Bernard et al. (2022) document the pivotal role of large buyers in determining supplier performance and market access. Using French customs data, Fontaine et al. (2020) show substantial variation in unit values across importers transacting with the same exporter, consistent with buyer-specific pricing. Similar patterns are found by Huang et al. (2021) in France, Chile, and China. Among the few studies to model buyer power directly, Morlacco (2019) estimates substantial oligopsony power among French importers, while Atkin et al. (2024) show that bargaining between Argentinian importers and foreign suppliers affects the price impact of trade policy.
Formal models of pricing in firm-to-firm trade remain limited. Dhyne et al. (2022) develops a network-based model of oligopoly in which suppliers’ markups increase with their relationship-specific market share. Our framework nests this model as a special case but extends it to incorporate bilateral bargaining and importer market power. Grossman et al. (2024) also study firm-to-firm pricing under Nash-in-Nash bargaining, emphasizing how supply-chain links adjust in response to trade shocks. In contrast, we take the trade network as fixed and focus on how bargaining and network rigidities interact to determine equilibrium prices and short-run pass-through.
We also contribute to the literature on the firm-level determinants of cost shock pass-through. A large body of work attributes incomplete pass-through to variable markups and strategic complementarities in price-setting (e.g., Amiti et al., 2014, 2019a). [^4] See also Berman et al. (2012), Auer and Schoenle (2016), and Garetto (2016) for firm-level evidence consistent with this class of models. We extend these insights to a bilateral oligopoly setting that reflects the structure of a large share of international trade. Our framework nests these existing mechanisms but allows for a richer characterization of both markup and cost channels, shaped by bargaining power and network position. Related work includes Gopinath and Itskhoki (2011) and Goldberg and Tille (2013), who model bargaining in firm-to-firm trade but abstract from importer market power and cost-side adjustment. Empirically, we find that these cost-based responses, rather than variable markups, are the primary driver of incomplete pass-through in these settings.
Finally, we contribute to the literature on shock propagation in production networks. A large body of work demonstrates that shocks can spread through input–output linkages and affect aggregate outcomes (Acemoglu et al., 2012; Di Giovanni et al., 2014; Grassi, 2018), often using natural experiments to trace transmission along supply chains (Barrot and Sauvagnat, 2016; Boehm et al., 2019; Carvalho et al., 2021). Recent work by Acemoglu and Tahbaz-Salehi (2025) highlights how bilateral bargaining can amplify supply chain disruptions through the price channel. We complement this perspective by modeling how idiosyncratic shocks to individual relationships affect prices under two-sided market power. While we abstract from general equilibrium feedback, our focus is on the within-network transmission of trade shocks through pricing, a key but underexplored margin in the propagation of shocks across production networks.
@@ 2 Theoretical Framework
This section develops a bargaining theory of firm-to-firm trade with two-sided market power. The model links markups and pass-through elasticities to importer and exporter market shares, along with a small set of structural parameters that capture key features of market structure upstream and downstream. The model yields analytic structural equations and testable predictions, which we later bring to the data using U.S. customs records.
We impose the following assumptions to keep the model tractable. First, we treat the trade network as fixed. This assumption captures the lock-in effects in firm-to-firm trade and allows us to focus on the price-setting problem while abstracting from firms’ decisions to form or sever links. [^5] This separation between extensive and intensive margin decisions is consistent with prior work showing that dynamic sourcing choices, while important for long-run outcomes, are not necessary to analyze the short-run price effects of trade shocks (Blaum et al., 2018). Second, we abstract from nominal rigidities, such as fixed-price contracts or currency denomination, as these are unlikely to materially affect tariff pass-through. [^6] Evidence from Amiti et al. (2020) supports this view, showing similar short- and long-run pass-through rates of 2018 U.S. import tariffs. See Goldberg and Tille (2013) for a theory of importer-exporter bargaining over the transaction price and exchange rate exposure. Third, we consider a static framework of single-product negotiations, despite our data reflecting repeated, multi-product interactions between firms. [^7] Repeated relationships affect contract structure and trade volumes (Macchiavello and Morjaria, 2015), exchange rate pass-through (Heise, 2024), and the propagation of shocks (Martin et al., 2023). We return to these features and discuss how we incorporate information on relationship duration in Section 3 and 4.

### v2

@@ Abstract
We develop and estimate a structural model of bargaining in firm-to-firm trade to study the determinants of tariff pass-through. The model features oligopoly and oligopsony power and yields analytical expressions for bilateral markups and pass-through based on two sufficient statistics: the supplier’s share in the buyer’s purchases and the buyer’s share in the supplier’s output. Using U.S. import data, we find substantial importer bargaining power and steep export supply curves. These primitives imply that cost changes, rather than markup adjustments, dominate pass-through, accounting for the bulk of incomplete pass-through of the 2018 U.S. tariffs and its heterogeneity across buyer-supplier links.
@@ 1 Introduction
International trade is organized around a network of buyers and suppliers that transact repeatedly across borders and jointly generate rents along global supply chains. How these rents are divided, and who bears cost shocks, has first-order implications for the incidence of trade policy. Yet canonical trade models abstract from bilateral relationships, thereby missing key mechanisms of shock transmission. We develop and estimate a structural model of firm-to-firm trade where two-sided market power determines the level and heterogeneity of relationship-level markups and pass-through. Applied to the 2018 U.S. tariffs, our model explains within-relationship pass-through as the outcome of bilateral bargaining between importers and exporters.
Pass-through of a cost shock reflects two distinct forces. The first is a markup channel, whereby the negotiated price-cost margin adjusts to the shock, reallocating surplus within the match. The second is a cost channel: at the higher effective price, large buyers contract demand, moving suppliers down their supply curves and lowering marginal cost. The relative strength of these forces is ultimately an empirical question that depends on underlying primitives. A central primitive is the inverse export supply elasticity, which not only governs oligopsony power but also shapes the balance between the two channels of transmission through the slope of exporter’s marginal cost curve. With bilateral bargaining, these elasticities are match-specific, generating systematic heterogeneity in pass-through across trading relationships that models with atomistic buyers or suppliers are unable to capture.
We estimate the model using pre 2017 U.S. import data and apply it to analyze the 2018 tariff episode. The estimates indicate strong importer bargaining power and steep export supply curves, consistent with substantial oligopsony power. At these parameter values, the cost channel drives tariff pass-through, with marginal cost adjustments dominating markup responses. Pass-through varies systematically across relationships and declines with buyer share, so larger importers face lower pass-through through a higher residual inverse supply elasticity. In the aggregate, pass-through is incomplete at 67 to 73 percent, implying that exporters absorb roughly one third of the tariff burden, through reductions in scale and quasi-rents. This match-level perspective contrasts with product-level studies that find near complete pass-through of the 2018 tariffs, [^1] We reconcile the match level and product level evidence later in the introduction. while reduced-form evidence supports the model predictions relative to alternatives without bargaining or upward-sloping export supply.
The model extends the network-based oligopoly framework of kikkawa2019imperfect to incorporate two-sided market power and decreasing returns to scale, which generate upward-sloping export supply and capture lock-in effects in global supply chains (antras2015global). Prices are set through bilateral Nash bargaining, with outside options given by each party’s alternative incumbent relationships. The equilibrium yields analytical expressions for bilateral markups and pass-through as functions of two sufficient statistics: the supplier share, measuring the exporter’s concentration among the buyer’s input expenditure, and the buyer share, measuring the importer’s concentration among the exporter’s total input supply.
Our framework offers a unified theory of exporter and importer market power in firm-to-firm trade. On the exporter side, rents arise from product differentiation and supplier concentration, expressed as an oligopoly markup over marginal cost that rises with the exporter’s supplier share, as in standard models (e.g. Atkeson2008). On the importer side, buyer power takes two distinct forms. Positive bargaining power gives a buyer countervailing power: the ability to extract a share of supplier rents, without affecting their scale. A positive inverse export supply elasticity further generates quasi-rents, resulting in an oligopsony markdown: a price-cost ratio below one that declines with the importer’s buyer share, as larger buyers capture a greater share of those quasi-rents through bargaining. [^2] In standard oligopsony models, oligopsony power operates through quantity distortions: buyers restrict demand to reduce the price they face. In our model, quantities always lie on buyers’ demand curves and there are no such distortions in equilibrium; the inverse supply elasticity nonetheless governs oligopsony power through the bargaining channel. See Section 2 for details. Under two-sided market power, the equilibrium price-cost margin is a convex combination of the oligopoly markup and the oligopsony markdown, with bargaining power as weight.
We study pass-through within existing buyer-supplier relationships: local price responses holding fixed all other match prices and general-equilibrium variables, and abstracting from longer-run adjustments driven by relationship entry and exit. A novel insight is that the two primitives that govern oligopsony power, namely the inverse export supply elasticity and bargaining power, also govern the pass-through determinants: the former determines whether pass-through reflects cost or markup adjustments; the latter shapes the nature of markup adjustment. When the inverse supply elasticity is high, demand contractions produce large movements in exporters’ marginal cost, and the cost channel drives most of the pass-through, regardless of bargaining power; when it is low, the cost channel is more muted and markup adjustments dominate. Within the markup channel, bargaining power governs the nature of strategic interaction: low buyer bargaining power implies strategic complementarities, which dampen pass-through (e.g., amiti2014importers; amiti2019international); high buyer bargaining power implies strategic substitutabilities arising from oligopsony power, which amplify it.
We take the model to U.S. transaction-level customs data covering the 2018 tariff episode. Our main data source is the U.S. Census Longitudinal Firm Trade Transaction Database (LFTTD), which records values and quantities of import transactions at the importer-exporter-product level. Our main sample focuses on repeated, arm’s-length relationships involving the exchange of intermediate and capital goods. Identification exploits cross-sectional price variation across buyers sourcing from the same exporter, product, and year: comparing prices across buyers within an exporter-product-year cell differences out exporter-level cost variation, so that remaining price differences reflect bilateral price-cost margins, which the model maps to observed market shares and unobserved parameters. We estimate importer bargaining weights of approximately 0.8, four times those of foreign suppliers, and a (short-run) returns-to-scale parameter around 0.5, implying an average residual inverse export supply elasticity of 0.3 for the average link. [^3] This estimate is below standard returns-to-scale estimates from the production function literature, reflecting a fundamentally different object: short-run supply constraints within existing relationships rather than long-run technological returns to scale. We discuss this in Section 4. Under constant returns to scale, this elasticity would be zero and both oligopsony power and the cost channel would vanish.
Price and pass-through patterns in the data closely align with the model’s predictions. Controlling for marginal costs, prices rise with the exporter’s supplier share and fall with the importer’s buyer share, consistent with two-sided market power. Out of sample, the model closely replicates the 2018 tariff episode, matching both the average level of pass-through (78% in the model, against 87% in the data) and its cross-sectional gradient: a one-unit increase in buyer share lowers pass-through by 0.42 in the model, against 0.41 in the data. Quantity responses point in the right direction but are estimated less precisely.
Specifications imposing constant returns to scale generate a near-zero interaction between pass-through and buyer share and are readily rejected. A specification retaining decreasing returns but setting bargaining power to zero performs comparably on this gradient: consistent with the theory, under strong decreasing returns the cost channel dominates and the gradient is insensitive to bargaining power. The two specifications differ, however, on average pass-through. Strong importer bargaining power implies strategic substitutabilities among buyers that partially offset the cost channel, while zero bargaining power implies strategic complementarities that reinforce it, pushing pass-through further below one. The baseline model outperforms its no bargaining counterpart in matching average pass-through (78% vs. 72%), as it lies closer to the empirical point estimate, although both fall within the confidence interval.
We use the estimated model to study aggregate pass-through in firm-to-firm trade and decompose tariff incidence within these relationships. Weighting by initial import shares, aggregate pass-through is substantially incomplete, with elasticities between 66 and 73%, below the unweighted average of 87% because larger (high-share) buyers experience the lowest pass-through. The adjustment operates almost entirely through the cost channel: shutting down markup elasticities leaves aggregate pass-through largely unchanged, whereas eliminating cost elasticities raises it significantly. Roughly one third of the tariff burden is thus borne by foreign exporters, absorbed not through compressed margins but through lost quasi-rents as lower demand moves exporters down their marginal-cost schedules.
A related literature documents near complete pass-through of the 2018 tariffs at the product level, implying that U.S. buyers bore most of the tariff burden (amiti2019impact; amiti2020s; flaaen2019production; fajgelbaum2020return; cavallo2020tariff). Our finding of incomplete pass-through within continuing relationships helps reconcile these results. Product-level estimates combine within-relationship price adjustment with changes in the set of active trading partners. Our results indicate that a full account of aggregate incidence requires understanding how the set of trading relationships responds to tariffs.
@@
Our paper contributes to three main strands of the literature.
First, the paper relates to the growing literature on importer market power. Early contributions emphasize the role of large buyers in shaping supplier outcomes (bernard2019production; bernard2022origins), while more recent firm-to-firm evidence documents systematic price discrimination across buyers, consistent with buyer power (Fontaine2019; huang2024firm; ignatenko2024market; macedoni2024pricing). Despite this evidence, formal theories of buyer power in international trade remain limited. [^4] Alternative explanations include endogenous matching, whereby more productive buyers source from a larger set of suppliers and intensify upstream competition (huang2024firm), and supplier-side price discrimination, under which buyers with stronger outside options obtain lower markups (ignatenko2024market). Exceptions include Morlacco2019, who estimates substantial oligopsony power among French importers, and atkin2024trade, who show that importer-supplier bargaining shapes the effects of trade policy.
We develop a tractable framework that nests kikkawa2019imperfect and extends it to incorporate importer oligopsony power via bilateral bargaining and upward-sloping export supply. These ingredients are quantitatively central: together they account for cross-sectional markup comovements and the level and heterogeneity of pass-through.
Second, the paper relates to the literature on determinants of incomplete pass-through using micro-level data. A large literature attributes incomplete pass-through to strategic markup adjustment under oligopoly (e.g., amiti2014importers; amiti2019international), while an alternative line emphasizes cost-side adjustment via nontraded distribution and retail costs in the importing country (goldberg2013structural). Our cost channel instead arises upstream, from exporters’ decreasing returns and demand contractions by large buyers. [^5] Prior work on pass-through under bargaining (gopinath2011search; goldberg2013bargaining) imposes constant marginal costs and therefore abstracts from oligopsony and the cost channel. Which channel prevails is state-dependent: in our setting, where short-run supply constraints bind, the cost channel dominates, while the standard markup approach is nested as a special case under constant marginal costs and applies naturally to longer-run horizons.
Finally, we contribute to the macro literature on shock propagation in production networks. A large body of work shows how input-output linkages affect the propagation of idiosyncratic shocks (Acemoglu2012; Grassi2018). [^6] Empirical evidence confirms these mechanisms. See, for example, barrot2016input, boehm2019input, and carvalho2021supply. A common feature of this literature is that prices are either taken as given or governed by fixed markups, so shocks propagate primarily through quantities. acemoglu2025macroeconomics allow for bilateral bargaining to study the consequences of supply chain disruptions. In their framework, however, price distortions play no role. We complement this approach by characterizing how unit prices respond to cost shocks under bilateral bargaining with decreasing returns, and showing that price adjustment itself is a quantitatively important transmission channel.
The remainder of the paper is organized as follows. Section 2 develops the theoretical framework and derives the results that guide the empirical analysis. Section 3 describes the data and presents reduced form evidence on prices and pass-through. Section 4 outlines the structural estimation strategy and reports the results. Section 5 examines aggregate pass-through in firm-to-firm trade. Section 6 concludes.
@@ 2 Theoretical Framework
This section develops a tractable pricing theory of firm-to-firm trade. We focus on how bilateral bargaining shapes prices and quantities within existing trade relationships, taking the network structure as given to isolate short-run tariff transmission. The framework is static and considers single-product negotiations; repeated interactions and multi-product relationships are incorporated in reduced form through the estimated bargaining parameters (see Section 4). We also abstract from nominal rigidities, such as fixed-price contracts or currency denomination, which appear to play a limited role in tariff pass-through. [^7] For example, amiti2020s shows similar short- and long-run pass-through for the 2018 U.S. tariffs. All proofs and derivations are in Appendix A.